OpenStudy (anonymous):

Could I get some help factorizing the following expression? -6x^2-16x+32

OpenStudy (waleed_imtiaz):

6x^2+16x-32=0 OKay.... NOw...... 6x^2+24x-8x-32=0 6x(x+4)-8(x+4)=0.......... So, x=-4 or x=4/3

OpenStudy (anonymous):

\[-2(3x^{2} +8x -16)\]\[-2(3x^{2}+ 12x - 4x -16)\]\[-2[(3x-4)(x+4)]\]

OpenStudy (anonymous):

Thank you. I still have some trouble understanding the first step though.. where does 24x-8x come from (waleed).. or 12x-4x (curiusshubham)

OpenStudy (waleed_imtiaz):

see @IHGR-PreIB we have to make the pairs whose sum or difference should equal to the middle term which consist power of x as one........ ANd these two pairs should also satisfy the product of 1st and last term co-efficients...... For example..... 6x^2+16x-32=0... Its like this.. Okay....... NOw the product of 1st and last term co-efficients =6(-32)=192..... so I made the pairs as 24 and 8... because when u multiply these..... 24*(-8)=-192.... and If U add and subtract them u will get 16x..... now U got it ?

OpenStudy (anonymous):

I think I do. Thx a lot

OpenStudy (waleed_imtiaz):

My pleasure.....!